OpenStudy (anonymous):

The x-intercept of a line is -5 and the y-intercept of the line is -2. What is the equation of the line? a. y=-5/2x-5 b. y=2/5x+2 c. y=5/2x-2 d. y=-2/5x-2

OpenStudy (anonymous):

1) the x- intercept is the point, at which y=0. a. y=-5/2x-5 y is zero as soon as "-5/2x" equal out the "-5". it is zero as soon as "-5/2x" = +5. when is "-5/2x" = 5 ?